A rock-climbing gym charges nonmembers $18 per day
to use the wall plus $7 per day for equipment rental. Members pay an
annual fee of $400 plus $5 per day for equipment rental. How many days
must both a member and a nonmember use the wall in one year so that
both pay the same amount?

Answer by stanbon(57245)

A rock-climbing gym charges nonmembers $18 per day
to use the wall plus $7 per day for equipment rental.
Cost = (18+7)d = 25d
----------------------------
Members pay an annual fee of $400 plus $5 per day for equipment rental.
Cost = 400 + 5d
-------------------------
How many days must both a member and a nonmember use the wall in one
year so that both pay the same amount?
----
Equation:
Solve:
25d = 400 + 5d
20d = 400
d = 20 days (Equal cost at 20 days)
=====================================
Cheers,
Stan H.
====================